The maximum amount of salt that can be dissolved in water at 100C is 391g/L. Calculate the molar concentration in this case, knowing that 1 mole of water has 18g and 1 mole of salt has 58.5g.
Answer #1

Given 391 gram salt can dissolve in 1L of water.

So, moles of salt that can dissolve = mass/molar mass

= 391/58.5 = 6.6837 gram

Now, molar concentration = moles of salt/volume of solution

= 6.6837/1.00 = 6.6837 M

If you have any questions please comment

If you satisfied with the solution please rate it thanks..